consider the following reaction :2Al + Cr2O3 = Al2O3 +2cr. what is the mass of Al2O3 that will be produced upon complete reaction of 1.20* 10^21 atoms of Al


1
Expert's answer
2021-12-15T10:17:34-0500

2Al + Cr2O3 = Al2O3 +2 Сr

1 mole of matter cntains 6.022×1023 elementary entities.

1.20x1021 atoms of Al equals to 1.20x1021/ 6.022×1023 = 0.2x10-2 moles.

According to the equation, n (Al2O3) = 1/2 x n (Al) = 1/2 x 0.2x10-2 = 0.001 moles.

n = m / M

m = n x M

M (Al2O3) = 102 g/mol

m (Al2O3) = n x M = 0.001 x 102 = 0.102 g